# Impact of 5G SA Logical Vulnerabilities on UAV Communications: Threat Models and Testbed Evaluation

### 04 Mar 2026 v1

@@ Abstract
This paper studies the impact of logical vulnerabilities in 5G Standalone networks on UAV command and control communication. The work considers three attacker positions in the architecture: a malicious UE connected to the same logical network as the UAV, an adversary with access to the 5G core, and a compromised gNodeB. To evaluate these scenarios, a testbed was built using Open5GS, UERANSIM, and Kubernetes. The environment reproduces a UAV–GCS communication system over a 5G SA network and allows controlled execution of attacks on different network interfaces. The experiments show that attacks at different architectural points can disrupt UAV operations, including manipulation of control commands and termination of data sessions. The results highlight the importance of isolation mechanisms in the 5G user plane and integrity protection in UAV command protocols.
@@ I Introduction
Unmanned aerial vehicles (UAVs) rely on continuous communication with the control station to execute commands and transmit telemetry. In scenarios connected by cellular networks, this communication becomes directly dependent on the functioning of the network infrastructure.
5G Standalone networks offer a communication environment with low latency and high transmission capacity. These characteristics allow the transport of control traffic and data between the UAV and the control station using the mobile network infrastructure itself.
Despite these properties, communication between applications and the network involves multiple architectural elements. Connected devices, network core functions, and access nodes participate in the transport and management of data sessions. Each of these points can influence communication behavior.
This work analyzes how attacks conducted at different positions in the 5G architecture can affect UAV command and control communications. Three threat models are considered: a malicious device on the same logical network, an adversary with access to the network core, and a compromised access node.
To evaluate these scenarios, a testbed based on Open5GS, UERANSIM, and Kubernetes was built. The environment replicates a 5G Standalone network and allows observation of the effects of attacks on communication between the UAV and the control station.
II Mobile Networks and 3GPP 5G
The 5G Standalone architecture defines a clear separation between control plane and user plane functions. At the core of the network are functions such as AMF, SMF, UPF, UDM, AUSF, and PCF, organized according to the Service Based Architecture model. Network access is achieved through gNodeB base stations, which connect user devices to the 5G core.
The control plane is responsible for mobility management, authentication, and session establishment. The user plane transports data between devices and the external data network through GTP-U tunnels. The creation and management of these tunnels is performed by the PFCP protocol between the SMF and UPF.
The architecture also introduces concepts such as network slicing and Data Network Name (DNN), which allow the creation of logical communication domains. Devices that establish sessions on the same DNN can share the same data domain depending on the isolation policies configured on the network. In certain scenarios, this configuration can allow direct communication between devices connected to the same slice.
These characteristics influence the system’s attack surface, as different points in the architecture can be exploited to affect communication between devices connected to the network.
III UAV–GCS Scenario in the Testbed
To evaluate attack scenarios, an experimental environment was developed that replicates a standalone 5G network in a controlled environment. The testbed was implemented using a Kind-based Kubernetes cluster, where the 5G core components were deployed in containers.
The network core uses the Open5GS implementation, including the AMF, SMF, and UPF functions. The access layer is simulated with UERANSIM, which provides gNodeB and UE device implementations. The environment uses a single PLMN configured with MCC 999 and MNC 70 and a single slice defined by SST 1 and SD 0x111111.
Three UE devices are used in the experimental scenario. The first represents the UAV, which runs a MAVLink simulator responsible for telemetry and control communication. The second represents the ground control station (GCS). The third device represents a malicious agent used in the experiments.
Depending on the threat model analyzed, the devices may be connected to the same gNodeB or to different gNodeBs. In specific scenarios, an additional pod is also created within the core network with access to SMF and UPF functions to simulate an internal adversary with access to the N4 interface.
This environment allows for the replay of attacks at different points in the 5G architecture and observation of their effects on communication between UAVs and GCS.
IV Threat Models and Attacks
@@
To analyze the impacts of logical attacks on the 5G network upon UAV-GCS communication, we considered three types of scenarios, divided as follows: (1) a malicious UE connected to the same network, targeting the UAV-GCS link; (2) an insider with access to the 5G network core (e.g., N4); and (3) a compromised gNodeB where the attacker fully controls the interface between the UEs and the core. For each of these threat models, the attacker’s capabilities, assumptions, and assets at risk will be described, along with an example attack in our testbed.
IV-A Threat Model 1: Rogue UE Impersonating the UAV
@@
As established by the 3GPP architecture, the notions of slice (S-NSSAI), DNN, and 5G VN Group are distinct concepts that are important for understanding this threat scenario. A slice is a logical partition of network functions and resources. A DNN identifies the data network to which a PDU session is established. In typical implementations, UEs that establish PDU sessions to the same DNN are connected to the same data network and may share the same addressing space, depending on the forwarding and isolation policies configured in the user plane. Thus, the 5G architecture by itself does not guarantee isolation between UEs that establish PDU Sessions to the same DNN. In this context, scenarios can arise in which multiple UEs share the same data domain, leaving it to the system operator to introduce additional configurations, as allowed by the 3GPP specification, for isolation and filtering between UEs or within the data network itself.
@@
For 5G LAN-type services, the 3GPP specification defines a 5G VN Group as a set of UEs that can communicate with each other and provides for unicast traffic forwarding between two UEs within the same 5G VN group. Thus, one UE can send traffic to another UE via the user plane (e.g., local switching at the UPF). Slice security is focused on isolation between slices.
@@
In a context where UEs are in the same slice (intra-slice), user devices can be regarded as accessible attack vectors, as they are outside organizational protection and may be operated by non-technical users. As discussed in, unauthorized access to a slice or its services can create an environment that is vulnerable to various attacks. Moreover, the very attachment to the slice can lead to privacy issues, for example by allowing correlation of users based on slice identification.
@@
Figure 1: Threat Model 1: UAV, GCS, and ROGUE UE attached to the same gNodeB.
To explore this scenario and demonstrate the impact on UAV operations, we implemented a C2 Injection attack at the application layer, also taking advantage of the fact that the main UAV applications that use MAVLink, such as ArduPilot, PX4, and QGroundControl, for example, do not use signing as a standard in their applications.
Since in this threat model the ROGUE UE can directly reach the legitimate UAV, because it is in the same slice and DNN, the attacker performs a network scan to identify the UAV (e.g., by probing port 14550), as seen in Listing  1. From this and targeting these vulnerabilities, the ROGUE UE can impersonate the control station by configuring the MAVLink system identifier as 255 (Ghost GCS), the default value associated with the GCS.
Listing 1: UDP scan for MAVLink port 14550 from ROGUE UE pod.
root@ueransim-rogue-6677bfc74d-7s2pm:/tmp# nmap -sU -p 14550 -Pn -n 10.45.0.2
Starting Nmap 7.95 at 2026-03-03 18:46 UTC
Nmap scan report for 10.45.0.2
Host is up.
PORT STATE SERVICE
14550/udp open|filtered unknown
Nmap done: 1 IP address (1 host up) scanned in 2.98 seconds
@@
In the first phase, considered reconnaissance and espionage, ROGUE attempts to establish a ”MAVLink” connection on UDP port 14550 towards the UAV, sending heartbeats impersonating the legitimate GCS and using the command MAV_CMD_SET_MESSAGE_INTERVAL to request the UAV to send the message GLOBAL_POSITION_INT, which returns its position and is necessary for the flooding algorithm to function. If successful, the attacker can obtain its real position (latitude and longitude) as well as other forms of telemetry, compromising the confidentiality of the UAV from the outset. Even if there is no response (for example, firewall or application restrictions), the algorithm still proceeds to the flooding phase.
@@
In the testbed, to simulate a real attack, the attacker directly accesses the Kubernetes pod of the ROGUE UE and executes the ‘nmap‘ application to scan the network and then execute the exploit on the identified port/IP. The script, as per Algorithm 1, sends the heartbeat and the telemetry request. The algorithm then waits a few seconds and starts sending the MAV_CMD_NAV_LAND commands.
The result observed in our tests in the testbed is that the UAV accepted the commands, forcing its eventual landing and end of the mission, regardless of the legitimate operator’s intentions(Listings 3). The failsafe based on the GCS heartbeat timeout does not activate, as the ROGUE also sends heartbeats as GCS. The central problem is the lack of authentication in the application, not the loss of the link.
@@
5:  Send(mastermaster, COMMAND_LONG(SET_MESSAGE_INTERVAL))
@@
Listing 2: Execution of the C2 injection exploit: BLIND attack with MAV_CMD_NAV_LAND commands
root@ueransim-rogue-66f7bfc74d-7s2pm:/tmp# /opt/rogue-venv/bin/python3 exploit.py 10.45.0.2
[*] TARGET SET: 10.45.0.2:14550
[*] Connecting as Secondary GCS (ID 255)...
[.] Waiting for telemetry data for 5 seconds...
[-] Drone ignored the data request (Possible UDP restriction).
[!] SWITCHING STRATEGY TO: BLIND ATTACK.
[!] The drone will be forced to land at current vertical position.
[*] Starting neutralization sequence (Mode: BLIND)...
[>] Firing command packet 0/50...
[>] Firing command packet 10/50...
[>] Firing command packet 20/50...
[>] Firing command packet 30/50...
[>] Firing command packet 40/50...
[SUCCESS] Commands sent. Check UAV status (Landing Initiated).
Listing 3: Threat Model 1: logs of the UAV state after the attack (forced landing).
... (repeated)
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] landed.
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
... (repeated)
IV-B Threat Model 2: Insider Threat
This threat model considers an insider adversary without radio equipment or UE registration, but with access to the 5G core network. In this scenario, it is assumed that the attacker reached the core (e.g., via compromised infrastructure, supply chain, or misconfigured isolation) and has access to both SMF and UPF, in particular, access to the N4 interface (PFCP, UDP port 8805), an environment already depicted in [1]. On this interface, SMF sends commands to UPF to create, modify, and delete PDU Sessions and the associated GTP-U tunnels [8]. In this context, messages can be manipulated to disrupt UE connectivity on the User Plane [1][6].

### 12 Mar 2026 v2

# Logical Attacks in 5G Standalone Networks: Impact on UAV C2 Communications

@@ Abstract
UAVs are increasingly deployed in critical applications and rely on 5G networks for long-range command-and-control (C2) connectivity. As the C2 channel is safety-critical, disruptions or manipulation of this communication channel may lead to loss of control, mission failure, or safety incidents. The architectural complexity of 5G standalone (SA) introduces logical attack surfaces that may affect such applications, yet the impact of logical vulnerabilities in the 5G architecture on UAV command-and-control carried over cellular infrastructure has received little attention. In this work, we develop a reproducible testbed that emulates 5G SA and integrates a UAV C2 channel using MAVLink over the 5G User Plane through Open5GS and UERANSIM. We define three threat models (rogue UE in the same slice and DNN, insider with access to the N4 interface, compromised gNodeB) and implement representative attacks. Our evaluation shows that a rogue UE can inject C2 commands and force the UAV to land; an insider can tear down PDU sessions via PFCP and trigger UAV failsafe; a compromised gNodeB can alter MAVLink navigation commands and redirect the UAV. The results demonstrate that logical attacks on the 5G architecture can compromise UAV C2 without breaking air-interface encryption, revealing cross-layer vulnerabilities between cellular infrastructure and UAV communication protocols. We provide a threat-model framework, experimental evidence, and mitigations (MAVLink signing, integrity protection on N3 and N4 interfaces) for operators and system designers deploying UAVs over 5G.
@@ I Introduction
Unmanned Aerial Vehicles (or UAVs) are increasingly used in critical applications such as infrastructure inspection, emergency response, logistics, and military operations, and increasingly rely on 5G networks for long-range command-and-control (C2) connectivity. UAV C2 communication is safety-critical: loss or manipulation of this communication channel can lead to loss of control, mission failure, or safety incidents [26], [12]. 3GPP has specified support for Unmanned Aircraft System (UAS) and C2 in 5G [1], [33]. However, the architectural complexity of 5G standalone (SA) introduces new logical attack surfaces across its Control Plane and User Plane interfaces surfaces that may disrupt such applications. Although several works have studied 5G security or drone security over direct links, little attention has been given to the effect of logical attacks on the 5G architecture targeting the UAV C2 over cellular infrastructure [34], [29].
In this work we develop a reproducible testbed that emulates 5G SA and integrates a UAV C2 channel. We define three threat models (rogue UE in the same slice/DNN, insider with N4 access and a compromised gNodeB), implement representative attacks, and evaluate their impact on telemetry and command delivery. Our contributions are: (i) a 5G UAV security testbed; (ii) a threat model framework for logical attacks on UAV C2 over 5G; (iii) experimental evaluation of C2 injection, PFCP session teardown, and navigation hijacking; (iv) analysis and mitigations (MAVLink signing, N4 and N3 protection) for cyber-physical systems over cellular.
The remainder of this paper is organized as follows. Section II presents background on 5G and UAV C2 and describes the testbed. Section III presents the threat models and experimental evaluation. Section IV reviews related work. Section V concludes.
II Background
This section summarizes the key concepts and system components used throughout the paper. We first outline the 5G standalone architecture, the separation of Control Plane and the User Plane, and the interfaces (N3, N4) and functions (AMF, SMF, UPF) relevant to the threat models considered in this work. We then describe the UAV–GCS C2 link and the MAVLink protocol used to transport C2 traffic over cellular networks. Finally, we present the 5G UAV testbed that emulates a 3GPP-aligned network and transports MAVLink between a UAV and a GCS over the 5G User Plane, which serves as the platform for evaluating the attacks described in Section III.
II-A Mobile Networks and 3GPP 5G
The standard 5G architecture (Figure 1), also called 5G standalone (SA), is specified by the 3rd Generation Partnership Project (3GPP). In this specification, the network is divided into two components: the Radio Access Network (RAN) and the 5G Core (5GC). User Equipment (UE) connects to the 5G network through the gNodeB, which is the radio base station of the 5G RAN. In turn, the gNodeB connects to the 5G core. The 5GC adopts what is called a Service-Based Architecture (SBA), where core functions expose services via HTTP and interact as producers and consumers. One of the main design choices of 5G is the separation between User Plane and Control Plane. The Control Plane carries signaling and session management, while the User Plane carries application data. This separation allows centralized control of session management while data traffic is forwarded in the User Plane [2], [3].
Within the 5GC’s SBA functions, three functions that are particularly relevant to this work: the Access and Mobility Management Function (AMF), Session Management Function (SMF), and User Plane Function (UPF). The AMF is responsible for UE registration, mobility, and connection management, terminates the UE’s Non-Access Stratum (NAS - the signaling layer between the UE and the 5GC), and participates in authentication and security context configuration. The SMF creates, modifies, and releases PDU Sessions and controls UPF behavior through the N4 interface. The UPF routes User Plane traffic between the RAN and external data networks, acting as the anchor for the PDU session traffic and the termination point for GTP-U tunnels from the gNB [2].
Figure 1: Overview of the 5G standalone architecture highlighting the Control Plane and the User Plane separation and the interfaces relevant to this work.
The 5G architecture separates user data and control messages. The Control Plane is used for control signaling between the UE and the 5GC (e.g., by N1 and N2) and between core functions (e.g., by service-based interfaces or N4). The User Plane carries application traffic from the UE and traverses the gNB and the UPF. The data flow follows the path UE → gNB → UPF → Data Network (DN) [2]. This separation is important because it allows the 5G architecture operator to manage sessions and policies in the Control Plane while routing data in the User Plane without passing through control nodes, for example. Among the main interfaces of these planes, primarily used for the research reported here, are interfaces N3 and N4. Interface N3 is located between the RAN and UPF and carries User Plane traffic. In N3, packets are encapsulated with GTP-U (GPRS Tunnelling Protocol for the User Plane), and the internal payload of the UE’s IP traffic [2]. Interface N4 is located between the SMF and UPF and uses the Packet Forwarding Control Protocol (PFCP) [4]. On interface N4, the SMF sends session establishment, modification, and removal requests to the UPF. The UPF applies the corresponding rules and creates or terminates the GTP-U tunnels used in N3. Thus, N4 controls which User Plane sessions exist and how they are routed.
UE connectivity within the 5G architecture is established through a PDU Session [2]. The UE requests a PDU Session for a Data Network Name (DNN), which identifies the external data network (e.g., the Internet or a private network). The SMF chooses a UPF, and the UPF allocates an IP address and configures the N3 tunnel towards gNodeB. Once the PDU Session is established, the UE → gNB → UPF → DN path is in effect. UEs using the same DNN and the same slice can share the same logical data network and address space, depending on the operator’s configurations.
One of the key features of the 5G architecture is support for network slicing [2]. A slice is a logical partition of network functions and resources, identified by an S-NSSAI (Slice/Service Type and Slice Differentiator). Multiple slices can coexist over the same physical infrastructure. The 3GPP specifications emphasize isolation between slices. However, isolation between UEs within a slice depends on the 5G network operator’s configuration [2].
Within the 5G architecture, as mentioned earlier, application traffic is transported in the User Plane. IP packets and their payloads (e.g., UDP or TCP and application protocols) are transported within the GTP-U tunnel in N3. However, User Plane protection is partly optional in 3GPP. Encryption and integrity at the radio interface (e.g., NEA2/NIA2) apply between the UE and the gNB [5]. gNodeB decrypts and decapsulates User Plane data to forward it to the N3 interface. Protection on N3 (e.g., IPsec) is optional and operator dependent.
II-B UAV Command and Control (C2)
An UAV is an aircraft operated without an onboard human pilot. The term UAS designates the set formed by the UAV, the human element, the command and control, the data communication link, and the payload and launch and recovery elements [12]. Similarly, the UAS can be described as the union of the aircraft (UAV), the data link, and the Remote Pilot Station (RPS), with the Ground Control Station (GCS) being the ground-based system that allows the pilot to command the drone and receive telemetry [29], [26].
UAVs rely on a C2 link to communicate with the GCS. This channel carries control commands in the uplink and telemetry/status information in the downlink (e.g., position, attitude, and system state). Loss or manipulation of this link can cause the pilot to lose control and lead to a forced landing, crash, or hijacking [26], [12]. Many systems define failsafe behaviors when the link is lost or corrupted, such as returning to a predefined point or initiating landing [12].
One application protocol used for C2 between the UAV and the GCS is the Micro Air Vehicle Link (MAVLink), considered a de facto standard in many autopilots and control stations [29], [18], [39]. MAVLink is a message-based protocol that uses checksums for message integrity verification at the protocol level [29], [25]. It supports control commands, telemetry, and status messages and is adopted in commonly used autopilots and GCS software [39], [25]. At the application layer, MAVLink does not require confidentiality protection and, unless signing is enabled, does not provide message authenticity or integrity against malicious modification. When deployed over mobile networks such as 5G, MAVLink traffic is protected on the radio link by mobile network security mechanisms and is typically transported as IP traffic (e.g., over UDP) in the User Plane [29], [2].
The MAVLink 2 protocol stands out for its support of message signing, aiming to accept only commands with a valid signature. The signing does not encrypt messages; it only defines which messages will actually be triggered [8]. The main problem is that this mode is not enabled by default in applications using MAVLink 2 and needs to be activated manually (e.g., in Mission Planner or in the autopilot parameters) [8], [6]. As a result, many deployments run the protocol without application-layer authentication, leaving the C2 link vulnerable to malicious actions.
Figure 2: UAV Communication Network (RF / Wi-Fi / 4G / 5G).
The C2 link can be carried over several communication media: proprietary radio links (with typical bands in 915 MHz, 2.4 GHz and 5.8 GHz for line-of-sight operation) and by Wi-Fi or cellular networks such as 4G and 5G [26], [12]. Cellular connectivity, especially 5G, is considered an option for operations known as Beyond Line of Sight (BVLOS), in which the drone operates without the pilot maintaining direct visual contact [34][29]. BVLOS requires stringent performance and dependability requirements for communications (e.g., end-to-end latency for C2 messages and reliability). 3GPP has specified support for UAS and C2 in 5G [1], [33]. Maintaining a reliable and secure C2 channel remains a challenge for drone operations and carrier-grade reliability for BVLOS, according to previous studies, is still far from being achieved in practice [33], [10].
II-C 5G UAV Experimental Testbed
The testbed used in the experiments emulates a 5G standalone (SA) network and the UAV–GCS command-and-control link over the User Plane, so that protocol-level attacks on the 5GC, RAN, or UEs can be evaluated in a controlled environment. The system under test (SUT) consists of the 5GC (AMF, SMF, UPF, and other SBA functions), a radio access node (gNodeB), and the UEs. The goal of the emulation is to reproduce a 5G architecture aligned with 3GPP specifications and to carry MAVLink C2 traffic between the UAV and the GCS over that path. Prior work has used Open5GS and UERANSIM to build 5G security testbeds [43][42]. We adopt the same base technologies and add the UAV/GCS roles and threat-model variants.
The 5GC is Open5GS [28], deployed on Kubernetes [14] via the Gradiant Helm charts [19], which provide the 5GC SBA functions. The RAN and UEs are implemented with UERANSIM [20]. The cluster is created with Kind [36], which runs Kubernetes nodes inside Docker [16] containers on a single host. Each UE pod receives a tun interface and an IP address from the UPF after PDU Session establishment. Over this interface, the UAV and GCS execute Python scripts [30] that implement a MAVLink-based C2 channel (as in Section II-B). The ROGUE UE pod can run custom attack or reconnaissance tools. The environment does not rely on a physical radio interface, as the gNodeB and UEs communicate over the cluster network, which keeps the scenario reproducible, realistic enough for protocol-level attack experiments.
The deployment is driven by a single bash script organized in numbered phases. An initial phase creates the Kind cluster and namespace and enables IP forwarding. The next phase installs the Open5GS Helm chart with a values file (e.g. ngc-values from the Gradiant 5G charts documentation), applies an override so that the UPF runs with the required privileges, waits for the core pods to be ready, and then hardens the AMF by removing the NIA0 and NEA0 algorithms from the AMF configuration, so that the radio link uses stronger algorithms. A following phase defines a ConfigMap with the gNodeB configuration (PLMN 999/70, NCI, TAC, slice S-NSSAI, AMF address). The final phases insert the UERANSIM elements (UAV, GCS, ROGUE, and other UEs and gNodeBs according to the threat scenario).
As observed in our experiments, running the 5GC and RAN inside Kubernetes yields several practical benefits. The cluster isolates core, gNodeB, and UEs into separate pods and services. Scaling or replacing a component only requires modifying manifests or Helm values. The same script and images produce the same topology on any host running Kind and kubectl, so that experiments are highly reproducible. Declarative configuration (YAML) makes it easy to add or remove UEs, change slice/DNN, or point a UE to a different gNodeB by editing ConfigMaps and Deployments. Threat models are switched by using different initialization scripts or by applying different gNodeB/UE layouts.
On the UAV and GCS pods, after the nr-ue process has established the PDU session and the tun interface is up, a shell script is executed that reads the tun IP and prompts for the peer IP (GCS IP on the UAV side, UAV IP on the GCS side). The script injects that IP into the Python MAVLink script and starts it. The UAV script listens for MAVLink on UDP port 14550 and sends to the GCS on port 14551. The implementation uses the pymavlink library [9] is used with MAVLink 2.0, which sends heartbeats and GLOBAL_POSITION_INT and handles COMMAND_LONG (e.g. arm, takeoff, land) and SET_POSITION_TARGET_LOCAL_NED. The GCS script sends to the UAV on 14550 and listens on 14551, with system ID 255 (GCS). Both scripts implement a simple failsafe mecanism: the UAV enters RTL if it does not receive a GCS heartbeat for a configured interval. The C2 channel is thus carried end-to-end over the 5G User Plane (UE $\rightarrow$ gNB $\rightarrow$ UPF $\rightarrow$ back to the other UE), so that any attack that drops or alters that path or injects traffic from the ROGUE UE affects the MAVLink stream in the same way as in a real deployment. The environment is controlled (fixed PLMN, slice, and DNN), reproducible (script from a clean cluster), and modular (UAV/GCS/ROGUE and gNodeB definitions can be swapped or duplicated for each threat model).
III Threat Models and Attacks
@@
To analyze the impacts of logical attacks on the 5G network upon UAV-GCS communication, we considered three types of scenarios, divided as follows: (1) a malicious UE connected to the same network, targeting the UAV-GCS link; (2) an insider with access to the 5G network core (e.g., N4); and (3) a compromised gNodeB where the attacker fully controls the interface between the UEs and the 5GC. For each of these threat models, the attacker’s capabilities, assumptions, and assets at risk will be described, along with an example attack in our testbed.
III-A Threat Model 1: Rogue UE Impersonating the UAV
@@
As established by the 3GPP architecture [2], [3], the notions of slice (S-NSSAI), DNN, and 5G VN Group are distinct concepts that are important for understanding this threat scenario. A slice is a logical partition of network functions and resources. A DNN identifies the data network to which a PDU session is established. In typical implementations, UEs that establish PDU sessions to the same DNN are connected to the same data network and may share the same addressing space, depending on the forwarding and isolation policies configured in the User Plane. Thus, the 5G architecture by itself does not guarantee isolation between UEs that establish PDU Sessions to the same DNN. In this context, scenarios can arise in which multiple UEs share the same data domain, leaving it to the system operator to introduce additional configurations, as allowed by the 3GPP specification, for isolation and filtering between UEs or within the data network itself.
@@
For 5G LAN-type services, the 3GPP specification [2] defines a 5G VN Group as a set of UEs that can communicate with each other and provides for unicast traffic forwarding between two UEs within the same 5G VN group. Thus, one UE can send traffic to another UE via the User Plane (e.g., local switching at the UPF). Slice security is focused on isolation between slices [3].
@@
In a context where UEs are in the same slice (intra-slice), user devices can be regarded as accessible attack vectors, as they are outside organizational protection and may be operated by non-technical users. Unauthorized access to a slice or its services can create an environment that is vulnerable to various attacks. Moreover, the very attachment to the slice can lead to privacy issues, for example by allowing correlation of users based on slice identification [27].
@@
Figure 3: Threat Model 1 - UAV, GCS, and ROGUE UE attached to the same gNodeB.
To explore this scenario and demonstrate the impact on UAV operations, we implemented a C2 Injection attack at the application layer, also taking advantage of the fact that the main UAV applications that use MAVLink, such as ArduPilot, PX4, and QGroundControl, for example, do not use signing as a standard in their applications [8], [6].
Since in this threat model the ROGUE UE can directly reach the legitimate UAV, because it is in the same slice and DNN, the attacker performs a network scan to identify the UAV (e.g., by probing port 14550). From this and targeting these vulnerabilities, the ROGUE UE can impersonate the control station by configuring the MAVLink system identifier as 255 (Ghost GCS), the default value associated with the GCS [25].
@@
In the first phase, considered reconnaissance and espionage, ROGUE attempts to establish a “MAVLink” connection on UDP port 14550 towards the UAV, sending heartbeats impersonating the legitimate GCS and using the command MAV_CMD_SET_MESSAGE_INTERVAL to request the UAV to send the message GLOBAL_POSITION_INT, which returns its position and is necessary for the flooding algorithm to function. If successful, the attacker can obtain its real position (latitude and longitude) as well as other forms of telemetry, compromising the confidentiality of the UAV from the outset. Even if there is no response (for example, firewall or application restrictions), the algorithm still proceeds to the flooding phase.
@@
Listing 1: Execution of the C2 injection exploit: BLIND attack with MAV_CMD_NAV_LAND commands.
[*] Connecting as Secondary GCS (ID 255)...
[.] Waiting for telemetry data for 5 seconds...
[-] Drone ignored the data request (Possible UDP restriction).
[!] SWITCHING STRATEGY TO: BLIND ATTACK.
[!] The drone will be forced to land at current vertical position.
[*] Starting neutralization sequence (Mode: BLIND)...
[>] Firing command packet 0/50...
[>] Firing command packet 10/50...
[>] Firing command packet 20/50...
[>] Firing command packet 30/50...
[>] Firing command packet 40/50...
[SUCCESS] Commands sent. Check UAV status (Landing Initiated).
In the testbed, to simulate a real attack, the attacker directly accesses the Kubernetes pod of the ROGUE UE and executes the nmap application to scan the network and then execute the exploit on the identified port/IP. The script, as per Algorithm 1, sends the heartbeat and the telemetry request. The algorithm then waits a few seconds and starts sending the MAV_CMD_NAV_LAND commands (Listing 1).
The result observed in our tests in the testbed is that the UAV accepted the commands, forcing its eventual landing and end of the mission, regardless of the legitimate operator’s intentions(Listings 2 and 3). The failsafe based on the GCS heartbeat timeout does not activate, as the ROGUE also sends heartbeats as GCS. The central problem is the lack of authentication in the application, not the loss of the link.
Listing 2: UAV log: repeated COMMAND_LONG (MAV_CMD_NAV_LAND, command=21) received from ROGUE.
... (repeated)
[UAV] received: COMMAND_LONG | content: {’mavpackettype’: ’COMMAND_LONG’, ’target_system’: 1, ’target_component’: 1, ’command’: 21, ’confirmation’: 0, ’param1’: 0.0, ’param2’: 0.0, ’param3’: 0.0, ’param4’: 0.0, ’param5’: 0.0, ’param6’: 0.0, ’param7’: 0.0}
[UAV] received: COMMAND_LONG | content: {’mavpackettype’: ’COMMAND_LONG’, ’target_system’: 1, ’target_component’: 1, ’command’: 21, ’confirmation’: 0, ’param1’: 0.0, ’param2’: 0.0, ’param3’: 0.0, ’param4’: 0.0, ’param5’: 0.0, ’param6’: 0.0, ’param7’: 0.0}
[UAV] received: COMMAND_LONG | content: {’mavpackettype’: ’COMMAND_LONG’, ’target_system’: 1, ’target_component’: 1, ’command’: 21, ’confirmation’: 0, ’param1’: 0.0, ’param2’: 0.0, ’param3’: 0.0, ’param4’: 0.0, ’param5’: 0.0, ’param6’: 0.0, ’param7’: 0.0}
[UAV] received: COMMAND_LONG | content: {’mavpackettype’: ’COMMAND_LONG’, ’target_system’: 1, ’target_component’: 1, ’command’: 21, ’confirmation’: 0, ’param1’: 0.0, ’param2’: 0.0, ’param3’: 0.0, ’param4’: 0.0, ’param5’: 0.0, ’param6’: 0.0, ’param7’: 0.0}
... (repeated)
Listing 3: logs of the UAV state after the attack (forced landing).
... (repeated)
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] landed.
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
... (repeated)
@@
5:  Send(mastermaster, COMMAND_LONG(MAV_CMD_SET_MESSAGE_INTERVAL, $msg\_id\leftarrow 33$, $interval\leftarrow 1000000$)) $\triangleright$ Request GLOBAL_POSITION_INT at 1 Hz (Phase 1: reconnaissance)
@@
III-B Threat Model 2: Insider Threat
This threat model considers an insider adversary without radio equipment or UE registration, but with access to the 5G core network. In this scenario, it is assumed that the attacker reached the core (e.g., via compromised infrastructure, supply chain, or misconfigured isolation) and has access to both SMF and UPF, in particular, access to the N4 interface (PFCP, UDP Port 8805), an environment already depicted in recent works [7]. On this interface, SMF sends commands to UPF to create, modify, and delete PDU Sessions and the associated GTP-U tunnels [4]. In this context, messages can be manipulated to disrupt UE connectivity on the User Plane [7][31].